# Efficient detection of chaos through the computation of the Generalized Alignment Index (GALI) by the multi-particle method

### v1

@@ 4.2 Hénon-Heiles model
Clearly large decay times close to the maximal time ($t=500$) are related to regular islands [yellow colored regions] while chaotic seas [purple colored regions] have in general smaller decay times ($t=100-200$). In addition, we also find initial conditions with intermediate values of decay times, which are in general at the edge of regular islands and belong to ‘sticky’ orbits. Although we are integrating more tangent vectors than the cases above, the advantage of this method is that we can stop the simulation for each set of tangent vectors once the GALI has reached the threshold. Following this approach, we obtain a computational time $t=28.6$ hours, close to what was obtained for the $\mbox{GALI}_{2}$, despite having to integrate an extra set of tangent vectors for $\mbox{GALI}_{3}$ than in the $\mbox{GALI}_{2}$ case. Thus, the results above not only confirm the ability of the GALI method implemented using the MPM to characterize the global dynamics of Hamiltonian systems, but also demonstrates its speed compared to the MLE approach.
Figure 10: $\bm{\beta-}$FPUT chain. The time evolution of the ftmLE $\lambda_{1}$ for two perturbations of SPO1 in the $\beta-$FPUT model with five particles: (a) A regular orbit with energy $\mathcal{H}_{5}=5$, and (b) a chaotic orbit with energy $\mathcal{H}_{5}=10$. Note that the solid lines correspond to results computed using the variational method (VM), while dotted and dashed lines show the results of the multi-particle method (MPM) with a deviation vector of norm $d_{0}=10^{-8}$ and $d_{0}=10^{-11}$ respectively. The dashed black line guides the eyes with slope $\approx-1$. In all these cases, apart from the deviation at the latest stage of the regular orbit evolution for the $d_{0}=10^{-11}$ vector, the multi-particle method accurately reproduces the variational method results.
@@ 4.3 The $\beta$-Fermi-Pasta-Ulam-Tsingou model
Figure 11: $\bm{\beta-}$FPUT chain. Time evolution of several computed GALIs for (a) and (b): A regular orbit of the 5-particle FPUT lattice, (c) and (d): A chaotic orbit. In all cases black lines correspond to computations using the variational equations, while colored lines show the results of the multi-particle method. In (a) and (c) a deviation vector of norm $10^{-8}$ was used, while for (b) and (d) a norm of $10^{-11}$ was used. The dashed grey lines show the theoretically predicted slopes for GALIk of regular orbits with $k>N$, according to Eq. (7). Note that in panels (a) and (b), time is plotted according to its logarithm, while panels (c) and (d) have a linear time scale.
@@
We now explore the performances of the GALI method using another well-known Hamiltonian model: the $\beta-$FPUT lattice [Eq. (39)]. We first look at a small lattice, consisting of five ($N=5$) particles, to establish the basic behaviours of the orbits we are going to study using the MLE. Knowing from the linear stability analysis that the value of the first energy destabilization threshold is $\mathcal{H}_{5}^{c}=7.4$ for the SPO1 with $N=5$, we choose as initial conditions (see B) perturbations of the SPO1s at fixed energy with $\mathcal{H}_{5}=5$ [Eq. (42)] and $\mathcal{H}_{5}=10$ [Eq. (43)] respectively below and above $\mathcal{H}_{5}^{c}$. For the integration of the equations of motion and variational equations, we fix integration time step $h=0.02$ leading to the relative energy error being always bounded from above by $E_{r}^{N}\approx 10^{-11}$. In addition, for both method we set the renormalization time $\tau=1$.
@@
As a representative case, the time evolution of the MLE for the initial condition with $\mathcal{H}_{5}=5$ is presented in Fig. 10, computed using the VM [continuous blue line] as well as the MPM. In the case of the MPM, we use the magnitude of the initial set of separation vectors $d_{0}=10^{-8}$ [dotted red line] and $d_{0}=10^{-11}$ [dashed green line]. As Fig. 10(a) shows, the MLE values decay towards 0 following the power law $t^{-1}$ (shown by the black dashed line) for the VM, confirming the regular nature of the orbit we are investigating. Using exactly the same deviation vectors as for the VM for the MPM calculations, with vectors rescaled to have norms $d_{0}=10^{-8}$ and $d_{0}=10^{-11}$, both accurately track the VM result until the last stage of the evolution where the results of the separation vectors with initial norm $d_{0}=10^{-11}$ starts yielding discrepant outcomes. On the other hand for the orbit with $\mathcal{H}_{5}=10$, we can see in Fig. 10(b) that all simulations show similar time evolution of the MLE, which converge to practically constant values of the MLE at the final time of the evolution. This clearly establishes chaotic nature of the orbit above. It is worth noting that the explanations for the inaccuracy of the MPM computations of the MLE with $d_{0}=10^{-11}$ can be found in Ref. [29], in particular relating to the limitations of working too close to machine precision.
@@
We investigate the same orbits as above using this time the GALI method. In Figs. 11(a) and (b) the black lines from the top to the bottom respectively depict the $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$, $\mbox{GALI}_{8}$ and $\mbox{GALI}_{10}$ calculated using the VM against time for the regular orbit with $\mathcal{H}_{5}=5$ of Fig. 10(a). In much the same way as for the HH model, the behaviour of the VM-calculated GALI values follow the theoretical predictions, irrespective of the choice of the initial size of the tangent vectors $d_{0}=10^{-8}$ [Fig. 11(a)] and $d_{0}=10^{-11}$ [Fig. 11(b)]. That is to say, we find that the time evolution of the $\mbox{GALI}_{2}$ and $\mbox{GALI}_{4}$ tend to saturate toward constant values, and the $\mbox{GALI}_{6}$ to $\mbox{GALI}_{10}$ decay following power laws against time up to the final time of our simulations. On the other hand, the same computations performed using the MPM for $\mbox{GALI}_{2}$ [blue curve], $\mbox{GALI}_{4}$ [orange curve], $\mbox{GALI}_{6}$ [green curve], $\mbox{GALI}_{8}$ [red curve] and $\mbox{GALI}_{10}$ [purple curve] show a less trivial picture. We find that the MPM-calculated GALI values practically coincide with the ones computed using the VM, for times up to $t\approx 10^{4}$ for $\mbox{GALI}_{10}$ and $\mbox{GALI}_{8}$, while this time is much larger than our computational time in case of $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$ and $\mbox{GALI}_{6}$ with $d_{0}=10^{-8}$, Fig. 11(a). However, this compatibility between the two methods lasts for a shorter time ($t\approx 10^{3}$) in the case of the computation using $d_{0}=10^{-11}$ in Fig. 11(b). Consequently, the deviation of the computed GALI values for growing time become more apparent in this context. Nonetheless, in both cases the computation of the $\mbox{GALI}_{2}$ remains rather robust since both methods return practically the same results for all simulations.
@@
When performing the same computations using the chaotic orbits with $\mathcal{H}_{5}=10$ [Fig. 10(b)], we expect all the GALI values to decay following a power law of time. This is exactly what is demonstrated by the time evolution of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$, $\mbox{GALI}_{8}$ and $\mbox{GALI}_{10}$ [black curves from top to bottom respectively in Fig. 11(c-d)] computed using the VM with $d_{0}=10^{-8}$ [Fig. 11(c)] and $d_{0}=10^{-11}$ [Fig. 11(d)]. In addition, the same calculation performed using the MPM tends again to follow similar temporal behaviors as that depicted by the VM for finite time $t\approx 50$. After this time, a clear deviation of the results of the MPM from those of the VM take place, now for all computed orders of the GALI.
Figure 12: $\bm{\beta-}$FPUT chain. Diagram of reliable regions for $\mbox{GALI}_{k}$ in the bFPUT computed by the multi-particle method for $N=5$, with the same chaotic initial conditions as used in Fig. 10. The total norm of the deviation vector is varied between $10^{-14}$ and $10^{-2}$. Panels (a)-(d) correspond to GALIs of order 2, 4, 6 and 8. The blue squares are the values of $\mbox{GALI}_{k}$ computed through the variational method, while the orange circles are the data computed via the multi-particle method. The GALI value is taken at a final time of $t=100$.
Let us now look for the region of reliability for GALI computed with the MPM, generated using the chaotic orbit of Fig. 10(b) and Figs. 12(c-d) with $N=5$ and $\mathcal{H}_{5}=10$. In our simulations we fix the integration time step $h=0.02$ and evolve the dynamical equations of the system up to $t\approx 200$ time units. Further, we set the renormalization time $\tau=1$, varying the $d_{0}$ values in the interval $[10^{-2},~{}10^{-14}]$, with unit step in logarithmic scale. The GALI values at the end of the integration time are recorded, and averaged over $100$ sets of random initial orthogonal tangent vectors.
In Fig. 12, the results of these computations are shown for $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$, and $\mbox{GALI}_{8}$ using the VM [blue squares] and MPM [orange dots]. Clearly the numerical uncertainties of $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$ and $\mbox{GALI}_{8}$ do not depend on $d_{0}$ for the VM as their values remain constant for all $d_{0}$ used [blue squares in Figs. 12(a-d) respectively]. On the other hand, for the MPM simulations whose results are represented by the orange filled circles, we observe an overall $V-$shape of the dependence of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$ and $\mbox{GALI}_{8}$ against $d_{0}$ with the minimum located at $d_{0}\approx 10^{-8}$ for all the displayed GALI. This is again in agreement with the prediction of Eq. (35).
@@
It is worth commenting a bit more on the $V-$shapes in Figs. 12(a-d). This $V-$shape appears to be more symmetric for small GALI orders [e.g. $\mbox{GALI}_{2}$ in Fig. 12(a)] compared to larger ones [e.g. $\mbox{GALI}_{8}$ in Fig. 12(d)] with respect to the vertical line passing through the minimum and perpendicular to the horizontal axis. The same tilted $V-$shape was also obtained with the HH model, increasing the time at which the reliable regions are computed for $\mbox{GALI}_{3}$ from $t\approx 200$ [Fig. 5(b)] to $t\approx 500$ and for $\mbox{GALI}_{4}$ from $t\approx 200$ [Fig. 5(c)] to $t\approx 300$. Note that symmetric $V-$shape of the reliable region of the GALI is the consequence of the leading order term approximation of the numerical error [Fig. 1] which seems to stand firm for the $\mbox{GALI}_{2}$. Consequently, the asymmetric $V-$shape of the reliable region of the $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$ and $\mbox{GALI}_{8}$ computed at time $t\approx 100$ in Figs. 12(b-d) respectively are indicative that the leading order approximation no longer holds for the accumulation of numerical errors. In fact, at large time lower order terms of the polynomial expansion of the numerical errors to the GALI also start contributing meaningfully to the global error.
In addition, we expect for even longer times the $V-$shape gives way to other asymmetrical effects, which may always possess local minimums at $d_{0}\approx 10^{-8}$. However, it is remarkable to see that even past the short-time accurate region of our analysis, the estimation of the reliable region remain accurate, i.e. $d_{0}\approx 10^{-8}$, which does significantly support the robustness of our findings.
It is worth emphasizing that we have found similar reliable region for the GALI method computed using the MPM with other chaotic orbits of the $\beta-\mbox{FPUT}$ models of various lattice sizes with $N=11$, $N=101$, and $N=501$ in Fig. 13 in C. These chaotic orbits are found in the neighborhood of unstable SPO1 with energy above the first energy destabilization with $\mathcal{H}_{11}^{c}=1.98$, $\mathcal{H}^{c}_{101}=1.515$ and $\mathcal{H}_{501}^{c}=1.502$ respectively for $N=11$, $101$ and $501$.
5 Conclusion and outlook
We have shown that it is possible to accurately detect chaos in multidimensional Hamiltonian systems with the Generalised Alignment Indices (GALI), using a multi-particle method (MPM) in place of the more common but complex technique involving the variational equations, which we call the variational method (VM). We performed a theoretical and numerical analysis of the errors associated with these MPM-GALI computations, and found that for short times this method is comparable in accuracy to the VM. The errors in the calculation of GALI depend primarily on the choice of deviation vector magnitude $d_{0}$, the renormalization time $\tau$, and the global truncation error of the numerical integration scheme $\varepsilon_{T}$ (Sec. 3). The latter being well approximated with the largest value reached by the relative energy error $E_{r}$ in case of Hamiltonian systems and for numerical integration schemes with fixed integration time step. Testing these predictions on two well-known Hamiltonian models, the Hénon-Heiles (HH) and $\beta$-Fermi-Pasta-Ulam-Tsingou ($\beta$-FPUT) systems, we confirmed in Sec. 4 that for sufficiently short times, which depend on the magnitude of the GALI, the MPM results match the VM results when a reasonable deviation vector magnitude is chosen. The choice of this magnitude is discussed and tested, and based on the numerical limit of double precision accuracy $\varepsilon=10^{-16}$, an initial deviation vector norm of $d_{0}\approx 10^{-8}$, a renormalization time $\tau\lesssim 1$ and a relative energy error $E_{r}\lesssim 10^{-8}$ are found to be optimal in all cases. The latter can be achieved easily using high order numerical integration schemes.
Beyond this, we also show that the MPM can be used to efficiently describe the global chaotic dynamics of the HH system, showing a clear functionality equivalent to that of the VM, and demonstrating that this technique is valid for the computation of GALIs in practice. Such high accuracy arises from the fact that short-time numerical simulations suffice to obtain a clear behavior of the GALI, in contrast with the MLE. This opens the way for GALI to be used in more complex systems, where the variational equations may be difficult/impossible to find, or simply add so much computational complexity as to be highly inefficient.
We believe this study complements existing work on the GALI method – all of which uses the VM – by enabling its application to a wider range of systems. Nevertheless, open questions remain such as considering the extension of these results to non-Hamiltonian systems, or considering short-time statistics of the GALI values for both regular and chaotic motion. This could ultimately lead to further improvements in the efficacy and speed of numerical computations with the GALI method.

### v2

@@ 4.2 The Hénon-Heiles model
To the question whether we can reach the same accuracy analyzing the phase space of the HH system using GALIs of higher orders, the answer is yes. First we remind the reader that the GALIs of higher orders decrease for both regular and chaotic orbits. Thus, we need to follow a different approach than recording the final value of the GALIs as was done for GALI2. In order to characterize the global dynamics of the HH system, we register the time it took the higher order GALIs to reach a particular small threshold value [18], exploiting the fact that the GALIs decay faster for chaotic orbits compared to regular ones. In Fig. 9(c), we color initial conditions according to the time it took $\mbox{GALI}_{3}$ to reach the threshold $10^{-8}$. This time is again averaged over $50$ different sets of initial orthogonal tangent vectors. Clearly large decay times close to the maximum considered time ($t=500$) are related to regular islands (yellow colored regions), while chaotic seas (purple colored regions) have in general smaller decay times ($t=100-200$). In addition, we also find initial conditions with intermediate values of decay times, which are in general at the edges of regular islands and belong to ‘sticky’ chaotic orbits. Although we are now integrating more tangent vectors than for obtaining the results of Figs. 9(a) and (b), the advantage of this approach is that we can stop the evolution of each orbit when the GALI3 has reached the threshold. Following this methodology, we required a computational time of $t=28.6$ CPU hours, close to what was obtained for the $\mbox{GALI}_{2}$ computations, despite having to integrate an additional tangent vector for $\mbox{GALI}_{3}$. Thus, all these results not only confirm the ability of the GALI method implemented using the MPM to characterize the global dynamics of Hamiltonian systems, but also demonstrate its computational speed and efficiency compared to the MLE approach.
@@ 4.3 The $\beta$-Fermi-Pasta-Ulam-Tsingou model
We now explore the performances of the GALI method using another well-known Hamiltonian model: the $\beta$-FPUT lattice [Eq. (39)]. We first look at a small lattice, consisting of five ($N=5$) oscillators, to establish the basic behaviours of the orbits we are going to study using the MLE. Knowing from the linear stability analysis performed in [57] that the value of the first energy destabilization threshold is $\mathcal{H}_{5}^{c}=7.4$ for the SPO1 with $N=5$, we choose (see B) orbits whose initial conditions are perturbations of the SPO1s at fixed energy with $\mathcal{H}_{5}=5$ [Eq. (42)] and $\mathcal{H}_{5}=10$ [Eq. (43)] respectively below and above $\mathcal{H}_{5}^{c}$. For the integration of the equations of motion and the related variational equations, we fix the integration time step to $h=0.02$, which results in the relative energy error (41) being always bounded from above by $E_{r}^{N}\approx 10^{-11}$. In addition, for both the VM and the MPM we set the renormalization time to $\tau=1$.
@@
As a representative case, the time evolution of the MLE for the regular with $\mathcal{H}_{5}=5$ is presented in Fig. 10(a), computed using the VM (continuous blue curve) as well as the MPM. In the case of the MPM, we use an initial separation vector of magnitude $d_{0}=10^{-8}$ (dotted red curve) and $d_{0}=10^{-11}$ (dashed green curve). As Fig. 10(a) shows, the MLE values decay towards zero following the power law $t^{-1}$ (shown by the black dashed line) for the VM, confirming the regular nature of the orbit we are studying. Using exactly the same separation vectors of the VM for the MPM calculations, rescaled to have norms $d_{0}=10^{-8}$ and $d_{0}=10^{-11}$, we see that we manage to accurately track the VM result until the last stage of the evolution when $d_{0}=10^{-8}$, while the results for $d_{0}=10^{-11}$ start deviating from the ones obtained by the VM for $t\gtrsim 10^{3.5}$. On the other hand, for the chaotic orbit with $\mathcal{H}_{5}=10$, we see in Fig. 10(b) that all simulations show similar time evolution of the MLE. In particular, the computed indices eventually converge to practically the same positive, constant value. This behavior clearly establishes the chaotic nature of the orbit. It is worth noting that the explanation for the inaccuracy of the MPM computations of the MLE for $d_{0}=10^{-11}$ can be found in Ref. [31], and is related to the limitations of the MPM when the norms of the separation vectors are too close to the machine precision.
Figure 10: $\bm{\beta}$-FPUT chain. The time evolution of the ftmLE $\lambda_{1}$ for two perturbations of the SPO1 of the $\beta$-FPUT model with $N=5$ oscillators: (a) A regular orbit for energy $\mathcal{H}_{5}=5$, and (b) a chaotic orbit with energy $\mathcal{H}_{5}=10$, whose initial conditions are given in Eq. (42) and Eq. (43) respectively. Note that in both panels the solid blue curves correspond to results computed using the VM, while red dotted and green dashed curves show results obtained by the MPM for an initial random deviation vector of norm $d_{0}=10^{-8}$ and $d_{0}=10^{-11}$ respectively. The dashed black line in (a) guides the eye for slope $-1$.
@@
We also study the behavior of the various GALIs for these two orbits of the $\beta$-FPUT model. In Figs. 11(a) and (b) the black curves from the top to the bottom respectively depict the $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$, $\mbox{GALI}_{8}$ and $\mbox{GALI}_{10}$ values calculated by the VM as a funtion of time $t$ of the regular orbit of Fig. 10(a). Similarly to the HH system, the behavior of the VM-calculated GALI values follow the theoretical predictions, irrespective of the choice of the initial norm of the tangent vectors, namely $d_{0}=10^{-8}$ [Fig. 11(a)] and $d_{0}=10^{-11}$ [Fig. 11(b)]. Thus, we see that the $\mbox{GALI}_{2}$ and $\mbox{GALI}_{4}$ values tend to saturate toward constant positive values, while the computed $\mbox{GALI}_{6}$, $\mbox{GALI}_{8}$ and $\mbox{GALI}_{10}$ decay following the power laws of Eq. (7) (which are depicted by dashed gray lines) until the final time of our simulations. On the other hand, the results obtained by the MPM for $\mbox{GALI}_{2}$ (blue curve), $\mbox{GALI}_{4}$ (orange curves), $\mbox{GALI}_{6}$ [(green curves), $\mbox{GALI}_{8}$ (red curves) and $\mbox{GALI}_{10}$ (purple curves) eventually show deviations from the theoretically expected behaviors. In particular, we find that the MPM-calculated GALI values $d_{0}=10^{-8}$ [Fig. 11(a)] practically coincide with the ones computed using the VM, for times up to $t\approx 10^{4}$ for the $\mbox{GALI}_{8}$ and $\mbox{GALI}_{10}$, while the results for the $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$ and $\mbox{GALI}_{6}$ are similar to the ones of the VM, until our final computation time. However, this compatibility between the two methods lasts for a shorter time ($t\approx 10^{3}$) in the case of the computation using $d_{0}=10^{-11}$ in Fig. 11(b), with only GALI2 results being similar to the VM ones until the end of our computations. It is worth noting that in both cases (i.e., for $d_{0}=10^{-8}$ and $d_{0}=10^{-11}$) the computation of the $\mbox{GALI}_{2}$ remains rather robust since both methods return practically the same results.
Figure 11: $\bm{\beta}$-FPUT chain. Time evolution of several computed GALIs for the regular [(a) and (b)] and the chaotic [(c) and (d)] orbits of Fig. 10. In all panels black curves correspond to results obtained by the VM, while colored curves show the results of the MPM. In (a) and (c) deviation vectors of norm $10^{-8}$ were used, while for (b) and (d) a norm of $10^{-11}$ was considered. The grey dashed lines in (a) and (b) show the theoretically predicted slopes for the various GALIk with $k>N$, according to Eq. (7). Note that in (a) and (b) time (the horizontal axis) is in logarithmic scales, while in (c) and (d) in linear scales.
@@
The GALI values of the chaotic orbit shown in Fig. 10(b) should decay to zero exponentially fast according to Eq. (8). This is exactly what is seen in time evolution of the $\mbox{GALI}_{2}$, $\mbox{GALI}_{4}$, $\mbox{GALI}_{6}$, $\mbox{GALI}_{8}$ and $\mbox{GALI}_{10}$ values [black curves from top to bottom respectively in Figs. 11(c) and (d)] computed using the VM with $d_{0}=10^{-8}$ [Fig. 11(c)] and $d_{0}=10^{-11}$ [Fig. 11(d)].On the other hand, the values obtained by the MPM tend again to follow similar temporal behaviors to the ones of the VM until $t\approx 80$ [Fig. 11(c)] and $t\approx 50$ [Fig. 11(d)]. After these times, a clear deviation of the results of the MPM from those of the VM is observed for all GALIs.
Let us now check the reliable region for the GALIs computed by the MPM, considering again the chaotic orbit of Fig. 10(b). In our simulations we fix the integration time step to $h=0.02$ and evolve the required orbits up to $t=200$ time units. We also set the renormalization time to $\tau=1$ and vary the $d_{0}$ values in the interval $[10^{-14},10^{-2}]$, with unit step in logarithmic scale. The GALI values at the end of the integration time are recorded, and averaged over $100$ sets of random initial orthogonal tangent vectors. The outcomes of this process are presented in Fig. 12, where results for both the VM (blue filled-triangles) and the MPM (red filled-circles) are shown for $\mbox{GALI}_{2}$ [Fig. 12(a)], $\mbox{GALI}_{4}$ [Fig. 12(b)], $\mbox{GALI}_{6}$ [Fig. 12(c)], and $\mbox{GALI}_{8}$ [Fig. 12(d)]. Clearly the numerical uncertainties of these GALIs do not depend on the value of $d_{0}$ for the VM as their values remain constant for all considered $d_{0}$. On the other hand, for the results of the MPM we observe an overall $V-$shape dependence of the GALI values against $d_{0}$ with the minimum located at $d_{0}\approx 10^{-8}$ for all cases. This behavior is again in agreement with the prediction of Eq. (35).
@@
It is worth commenting a bit more on the $V-$shapes of Fig. 12, which appear to be more symmetric for small GALI orders [e.g. $\mbox{GALI}_{2}$ in Fig. 12(a)] compared to larger ones [e.g. $\mbox{GALI}_{8}$ in Fig. 12(d)] with respect to the vertical line passing through the minimum. Although we do not show these results, in order to avoid repetitions of similarly looking plots, analogous, slightly tilted $V-$shapes were also obtained for the HH model as the time at which results were computed increased from $t=200$ [Fig. 5(b)] to $t=500$ for $\mbox{GALI}_{3}$, and from $t=200$ [Fig. 5(c)] to $t=300$ for $\mbox{GALI}_{4}$. We emphasize that the symmetric nature of the $V-$shape of the reliable region of the GALI values is a consequence of the leading order term approximation of the numerical error [as was graphically represented in Fig. 1]. This property holds true for the $\mbox{GALI}_{2}$ computations. Consequently, the asymmetric $V-$shape in the results for the $\mbox{GALI}_{4}$ [Fig. 12(b)], $\mbox{GALI}_{6}$ [Fig. 12(c)] and $\mbox{GALI}_{8}$ [Fig. 12(d)] at time $t\approx 200$ indicates that the leading order approximation of the error is no longer very accurate. In fact, at large times higher order terms of the polynomial expansion of the numerical errors of the GALI computation should also start contributing meaningfully to the global error. Nevertheless, it is remarkable to see that even in cases beyond the validity of our first order approximation, our estimation about the optimal $d_{0}$ value (i.e., $d_{0}\approx 10^{-8}$) remains accurate, something which strongly supports the robustness of our findings. It is worth emphasizing that we have found similarly looking reliable regions for the GALI method computed using the MPM for other chaotic orbits of $\beta$-FPUT models with various lattice sizes. Some of these results, in particular for $N=11$, $N=101$, and $N=501$ are shown in Fig. 13 of C.
5 Conclusions and outlook
We have shown that it is possible to accurately detect chaos in multidimensional Hamiltonian systems by the GALI technique, using the MPM in place of the more commonly applied VM, which involves the system’s variational equations that require the related equations of motion to be continuous and differentiable. We performed both a theoretical and a numerical analysis of the errors associated with the application of the MPM to evaluate the GALIs, and found that for relatively short evolution times the method’s accuracy is comparable to that of the VM.
The numerical errors in the calculation of the GALI when utilizing the MPM depend primarily on the choice of the deviation vector magnitude $d_{0}$, the renormalization time $\tau$, and the global truncation error of the numerical integration scheme $\varepsilon_{T}$ (Sect. 3). The latter error can be estimated by the largest value of the relative energy error $E_{r}$ in the case of Hamiltonian systems.
Testing these theoretical predictions for two well-known Hamiltonian models, namely the 2D HH system and the $\beta$-FPUT model for various values of its degrees of freedom, we confirmed in Sect. 4 that, for sufficiently short times, whose duration depend on the order of the GALI (i.e., a higher GALI order results to smaller times), the MPM results match the ones obtained by the VM when a reasonable deviation vector magnitude is chosen. The choice of this magnitude was discussed and investigated in detail, along with its relation to the numerical limit of double precision accuracy ($\varepsilon\approx 10^{-16}$). We found that an initial deviation vector norm of $d_{0}\approx 10^{-8}$, a renormalization time $\tau\lesssim 1$ and a relative energy error $E_{r}\lesssim 10^{-8}$ are optimal in all cases. Furthermore, we demonstrated that the MPM can be implemented to efficiently capture the global chaotic dynamics of the HH system, providing functionally equivalent results to the VM, something which constitutes a further confirmation of its validity for the practical computation of the GALI.
It is worth noting that the introduced MPM extends the idea of computing the most commonly used chaos indicator (i.e., the MLE) by considering two nearby orbits instead of solving the system’s variational equations to the case of the GALI chaos detection method. This is a computationally much harder task, as the evaluation the GALI of order $k$ by the MPM requires the simultaneous evolution of $k$ orbits in the neighborhood of the tested trajectory (along with the tested orbit itself), while for the estimation of the MLE only one additional orbit is needed. Nevertheless, the GALI method is known to have many advantages over the computation of the MLE [17, 60, 18]. Thus, our work provides a way to exploit these advantages even when the VM cannot be used.
In the framework of the MPM, it will be interesting in the future to investigate the possible ability of the short-term statistics of GALI values to enhance the efficiency and speed of numerical computations allowing the early discrimination between regular and chaotic motion, a feature which could possibly facilitate the integration of this approach with neural network systems [64].
Our work paves the way for the exploration of chaotic dynamics using the GALI chaos detection method in a broader range of dynamical systems, including those where variational equations are difficult, or even impossible to derive analytically, or when their formulation introduces significant computational complexity. Such dynamical systems could be models whose equations of motion exhibit discontinuities, or systems having empirical, non-Newtonian, non-Hamiltonian and higher-dimensional field vectors. For instance, extending our investigation of the applicability and performance of the MPM to compute GALIs for the HH system, it would be interesting to study the characteristics chaos for the relativistic version of the system [65, 66] using similar approaches. Another interesting direction for applying the methodology developed in our work is to study chaos in pattern forming and turbulent far-from-equilibrium systems. Such studies could include active fluids and solids whose equations of motion often extend the classical form including time-dependent or asymmetric stress/strain tensors. Investigations of this kind could be beneficial in novel applications like in swarm and soft robotic designs and bioengineering.
Figure 12: $\bm{\beta}$-FPUT chain. Reliable regions for the computation of the (a) GALI2, (b) GALI4, (c)GALI6, and (d) GALI8 of the chaotic orbit of the $\beta$-FPUT lattice considered in Fig. 10. using the VM (blue triangles) and the MPM (red circles). The presented GALI values have been obtained by averaging over $100$ sets of initial orthogonal tangent vectors. The final time of integration is $t=200$ time units.